# Geometry-Informed Neural Operator Transformer

### v4

@@ Abstract
Machine-learning-based surrogate models offer significant computational efficiency and faster simulations compared to traditional numerical methods, especially for problems requiring repeated evaluations of partial differential equations. This work introduces the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions on arbitrary geometries. GINOT employs a sampling and grouping strategy together with an attention mechanism to encode surface point clouds that are unordered, exhibit non-uniform point densities, and contain varying numbers of points for different geometries. The geometry information is seamlessly integrated with query points in the solution decoder through the attention mechanism. The performance of GINOT is validated on multiple challenging datasets, showcasing its high accuracy and strong generalization capabilities for complex and arbitrary 2D and 3D geometries.
@@
GINOT effectively processes complex, arbitrary geometries and varying input conditions with high predictive accuracy
@@ 1 Introduction
To address these challenges, data-driven approaches, often referred to as surrogate models, have been developed. These methods utilize deep learning techniques to approximate PDE solutions, offering significant reductions in computational costs and faster simulations without compromising accuracy. Such advancements are particularly advantageous for applications requiring repeated PDE evaluations [1], including sensitivity analysis [2, 3], gradient-based optimization [4, 5], uncertainty quantification [6, 7], and inverse design enabled by generative AI [8, 9].
@@
where ${G}_{\theta}$ is the neural operator with learnable parameter $\theta$, $\mathcal{F}$ is the space of input functions, and $\mathcal{G}$ is the space of output functions on query points x. Examples such as the Fourier Neural Operator (FNO) [10] and the Deep Operator Network (DeepONet) [11] have shown significant success in this area [12, 13, 14, 15, 16, 17, 18, 19, 20]. These models map inputs like initial and boundary conditions, material properties, and other functions to solution fields at query points. However, their applicability is often limited to simple geometries. For instance, FNO is tailored for fixed, regular meshes such as image grids or 3D voxels and requires a complete mesh graph as input due to its reliance on the Fast Fourier Transform kernel, making it unsuitable for arbitrary geometries and less scalable. While DeepONet can predict solution fields at arbitrary query points, its original formulation is restricted to single geometries with fixed query point locations and counts.
@@
However, real-world problems frequently involve complex and arbitrary computational geometries. Recent advances, including graph neural operators (GNO) [21], Geo-FNO [22], GI-FNO [23], and Geom-DeepONet [24], have addressed challenges in handling diverse geometries. GNO can process irregular grids using graph representations but is limited to local operations under constrained computational budgets, resulting in lower accuracy. Both Geo-FNO and GI-FNO rely on projecting irregular meshes onto fixed regular grids and mapping them back. However, Geo-FNO requires a fixed number of query points during training, while GI-FNO constructs new graphs on regular meshes using local kernel integration layers through GNO, which need longer inference time. Geom-DeepONet, while overcoming the limitation of single geometries of DeepONet by removing or repeating nodes to ensure fixed data size, is restricted to parametric geometries (e.g., length, thickness, radius) and does not generalize to fully arbitrary shapes. Additionally, GI-FNO requires a signed distance function (SDF) as input to represent geometry, and Geom-DeepONet depends on both geometry parameters and the SDF. However, obtaining the SDF is not always easily, particularly for complex 3D geometries represented by surface meshes, as it incurs a computational cost of $O(N_{g}\times N_{m})$ for distance calculation and $O(N_{g}\times N_{m})$ for the minimum value, where $N_{g}$ is the number of grid points in the domain and $N_{m}$ is the number of surface mesh—both of which are large for intricate geometries.
@@
Researchers have leveraged the attention mechanism of transformers [25] in neural operators to efficiently integrate information from input functions and query points [26, 27, 28, 29]. Neural Operator Transformers (NOT) are particularly advantageous for irregular meshes due to their attention mechanism, enabling the prediction of solution fields at arbitrary query points. Unlike DeepONet, which employs a simple dot product to combine information from input functions and query points, the attention mechanism in NOT directs each query point to focus on the most relevant information from the input functions. This capability allows NOT to generalize effectively and capture more complex relationships between input functions and query points. Building on this, the authors developed an SDF-based NOT that accommodates arbitrary geometries [9]. The SDF representation facilitates handling arbitrary geometries on fixed, regular grids, which are compatible with existing deep learning models. However, as noted earlier, obtaining the SDF is not always straightforward or computationally easily.
@@
In this work, we introduce the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries without requiring SDF as geometry features. GINOT consists of two primary components: the geometry encoder and the solution decoder. Drawing inspiration from Shap-E [30] by OpenAI, we design the geometry encoder to process the boundary point cloud, representing the geometry, into the KEY and VALUE matrices for the attention mechanism. This choice facilitates the integration of information from query points in the solution decoder.
@@
The geometry encoder utilizes sampling and grouping layers [31, 32] to extract local geometric features from the boundary point cloud. These local features are fused with global geometry information through cross-attention layers, enabling effective processing of unordered and non-uniformly distributed point clouds—unlike convolutional neural networks (CNNs), which operate on regular grids with uniform density. Because the number of points in the point cloud can vary between geometries, padding is applied to standardize input sizes within a batch. To ensure that padding points do not influence the geometry encoder’s output, a masking mechanism is employed: it prevents padding points from being selected in the sampling and grouping layers and excludes them from contributing to attention scores in the cross-attention layer.
@@ 2 GINOT
GINOT integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries. It consists of two primary components: the geometry encoder and the solution decoder, as illustrated in Fig. 1. Both components leverage the attention mechanism inherent in the transformer architecture [25], which facilitates the model’s ability to focus on relevant input information. The attention mechanism is defined as
@@
where $Q\in\mathbb{R}^{N_{q}\times d_{e}}$ represents the QUERY matrix, and $K,V\in\mathbb{R}^{N_{k}\times d_{e}}$ are the KEY and VALUE matrices, respectively. Here, $d_{e}$ denotes the embedding dimension, $N_{q}$ is the length of the query sequence, and $N_{k}$ is the length of the KEY and VALUE sequences. To ensure stable training, residual connections and layer normalization are applied within each attention block, as depicted in the bottom right of Fig. 1.
@@ 2.1 Geometry encoder
To address these challenges, we developed a geometry encoder that processes the boundary point cloud of a geometry into the KEY and VALUE matrices for the attention mechanism. This enables seamless integration of geometry information with query points in the solution decoder. In contrast to CNN inputs, boundary point cloud used to represent geometries have three key characteristics: (1) invariance to the order of points, (2) non-uniform point density across locations, and (3) a variable number of points across geometries. Drawing inspiration from Shap-E [30] by OpenAI, the geometry encoder, as shown in Fig. 1(a), is designed to handle these characteristics. It ensures invariance to point order and padding while maintaining robustness to variations in point density.
@@
The point cloud is first processed through a sampling layer and a grouping layer for local geometry features, originally introduced in PointNet [31] and PointNet++ [32], which are well-suited to handle the aforementioned properties. The sampling layer employs the iterative farthest point sampling (FPS) method [32] to select $N_{s}$ points from the point cloud. This method begins by randomly selecting a point, then iteratively selects the farthest point from the already selected points until $N_{s}$ points are chosen. Compared to random sampling, the FPS method provides better coverage of the entire point set for a given number of centroids. The selected points are then passed to the grouping layer, which forms $N_{s}$ groups, each containing $N_{p}$ points.
@@
For each group, the $N_{p}$ points are selected as those within a ball of radius $r$ (a hyperparameter) centered at the corresponding sampled point. If fewer than $N_{p}$ points exist within the ball, the group is padded with the nearest points to the center. Conversely, if more than $N_{p}$ points are present, the nearest $N_{p}$ points are chosen. Consequently, the output of the grouping layer has dimensions $N_{s}\times N_{p}\times d$, where $d$ represents the dimensionality of the point cloud. The sampling and grouping layers allow to extract local geometry features from nonuniform density point cloud [32]. These layers output the selected points along with their indices in the original point cloud, which are subsequently used in downstream layers.
@@
The point cloud is also processed through a positional encoder layer utilizing the Nerf encoding method [33], followed by a linear layer. Using the indices from the sampling and grouping layers, the output is indexed to extract local geometry information, resulting in a tensor of shape $N_{s}\times N_{p}\times C$, where $C$ represents the output channel dimension of the positional encoder. This tensor is concatenated with the output of the grouping layer. The combined tensor is then passed through a series of 2D convolutional layers and a multi-layer perceptron (MLP), which aggregates the information within each group, reducing the dimension of $N_{p}$ to 1. As a result, the final output has a shape of $N_{s}\times C$.
@@
where the output is unaffected by permutations of the index $m$, which is a dummy index under Einstein summation notation.
The sampling and grouping layers, combined with the attention mechanism, allow that the geometry encoder handle the point cloud with varying points density, while remaining robust to the permutation of points order. Since the number of points in the point cloud varies across different geometries, padding is typically required to standardize input sizes within a batch for deep learning models. To prevent padding points from influencing the geometry encoder’s output, a masking mechanism is employed. This mechanism excludes padding points from being selected in the sampling and grouping layers. Additionally, padding points can interfere with attention scores in the cross-attention layer of the geometry encoder. To mitigate this, the corresponding entries in the attention score matrix $(Q\cdot K^{T})/\sqrt{d_{e}}$ are set to "$\infty$- ∞", ensuring that the attention layer’s output is unaffected by padding points. The flexibility and robustness of the proposed geometry encoder are illustrated in experiments Section 3.
@@ 2.2 Solution decoder
The solution decoder predicts the solution field at the query points using the geometry information from the geometry encoder. First, the query points are processed through a Nerf positional encoding, followed by an MLP, to generate the QUERY matrix with dimensions $N_{q}\times d_{e}$, where $N_{q}$ is the number of query points and $d_{e}$ is the embedding dimension. This QUERY matrix is then passed into cross-attention layers, which integrate the geometry information from the geometry encoder. The attention mechanism ensures that the query points focus on the most relevant geometry features. Finally, the output of the cross-attention layers is decoded into the solution field at the query points using another MLP. The architecture of the solution decoder is illustrated in Fig. 1(b).
@@ 2.3 Extending GINOT
Many problems involve not only varying geometries but also additional input parameters, such as loading conditions, material properties, or boundary conditions. To address such cases, GINOT can be extended by incorporating additional encoders to process these input functions. The outputs of these additional encoders are concatenated with the output of the geometry encoder and passed through an MLP for aggregation of all information from the geometries and the additional inputs. The aggregated output is then used as the KEY and VALUE in the cross-attention blocks of the solution decoder. The architecture of the extended GINOT is illustrated in Fig. 2. This extension is demonstrated in the bracket lug example in Section 3.4, where varying loading conditions are included as additional inputs.
@@ 3 Experiments
where $y^{\text{true}}$ and $y^{\text{pred}}$ represent the ground truth and predicted solutions, respectively. For some datasets, the $L_{2}$ relative error may not be an ideal metric. Therefore, we also employ the normalized root mean square error (NRMSE) as an additional metric defined as
$\text{NRMSE}=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y^{\text{true}}_{i}-y^{\text{pred}}_{i})^{2}}}{y^{\text{true}}_{\text{max}}-y^{\text{true}}_{\text{min}}}.$ (5)
@@ 3.1 Overview
To evaluate the performance and accuracy of GINOT, six datasets are utilized: elasticity [22], Poisson equation with both structured and unstructured meshes, bracket lugs [24], micro-periodic unit cell (PUC) [34] in [9], and Jet Engine Bracket (JEB) [35]. Table 1 provides a detailed summary of these datasets, including their size, the number of query points, the number of points in the boundary point cloud, whether the dataset is parametric or fully arbitrary, and the type of solution considered. The first three datasets have a constant number of points in the boundary point cloud, while the last three datasets feature a varying number of points in the boundary point cloud. All datasets feature non-uniform point densities, as illustrated in the visualization results in the following subsections.
@@
DataSet Elasticity[22] Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug [24] Micro- PUC [34] JEB [35]
@@
Number of query point 972 2737 [1566,2838] [5054,15210] [1136,6363] [18085,55009]
Number points in boundary PC 45 144 144 [2184,4284] [180,484] [2253,10388]
@@
Solution Stress Mises stress Mises stress displacements Mises stress
Table 2 summarizes the training hyperparameters of the GINOT models. For the micro PUC dataset, two models are trained: one for predicting the solution at the final strain step and another for predicting the solution over historical strain steps.
@@
Elasticity Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug Micro PUC (Final strain step) Micro PUC (historical strain steps) JEB
Batch size 20 32 32 32 64 64 32
Optimizer Adam Adam Adam Adam Adam Adam Adam
Initial learning rate 0.001 0.001 0.001 0.001 0.001 0.001 0.001
@@
Grouping $r$ 0.2 0.2 0.2 0.5 0.2 0.2 0.5
Att. heads (Decoder) 4 8 8 8 4 4 1
@@
Table 3 summarizes the computational efficiency of GINOT. Training tasks were conducted on single NVIDIA A100 GPU on the DELTA machine or single NVIDIA H100 GPU on the DeltaAI machine, both hosted at the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign (UIUC). Inference tasks were executed on single NVIDIA A100 GPU on the DELTA machine.
@@
Elasticity Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug Micro PUC (Final strain step) Micro PUC (historical strain steps) JEB
@@
Table 4 summarizes GINOT’s performance across the datasets, presenting the $L_{2}$ relative error or NRMSE for both training and testing samples. These results are compared against baseline methods, with the best performance highlighted in bold. The findings demonstrate GINOT’s ability to tackle more complex problems while outperforming the baseline methods. Detailed experimental results are provided in the subsequent subsections.
@@
Dataset, Metric Data Geo-FNO[22] Geom- DeepONet[24] GNO[21] GI-FNO[23] GINOT
Elasticity, $L_{2}$ Training 1.25% NA 11.3% 1.32% 0.936%
Testing 2.29% NA 12.3% 1.66% 1.33%
Poisson, $L_{2}$ (structure mesh) Training 0.43% NA 7.43% 0.42% 0.39%
@@
Poisson, $L_{2}$ (unstructure mesh) Training NA 7.98% 0.49% 0.41%
@@
Bracket lug,$L_{2}$ Training 1.818% 10.5% 2.51% 0.378%
Testing 1.824% 11.3% 2.81% 0.45%
Micro PUC, $L_{2}$ (Final strain step) Training 16.5% 9.24% 7.93%
Testing 18.2% 9.95% 9.05%
Micro PUC, $L_{2}$ (historical strain steps) Training 18.9% 9.56% 8.42%

### v5

@@ Abstract
Machine-learning-based surrogate models offer significant computational efficiency and faster simulations compared to traditional numerical methods, especially for problems requiring repeated evaluations of partial differential equations. This work introduces the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions on arbitrary geometries. GINOT employs a sampling and grouping strategy together with an attention mechanism to encode surface point clouds that are unordered, exhibit non-uniform point densities, and contain varying numbers of points for different geometries. The geometry information is seamlessly integrated with query points in the solution decoder through the attention mechanism. The performance of GINOT is validated on multiple challenging datasets, showcasing its accuracy and generalization capabilities for complex and arbitrary 2D and 3D geometries.
@@
GINOT effectively processes complex, arbitrary geometries and varying input conditions with good predictive accuracy
@@ 1 Introduction
Data-driven surrogate models have emerged to alleviate these computational burdens. Leveraging deep neural architectures, they approximate PDE solution operators with substantial speedups while retaining acceptable accuracy, making them effective when many evaluations are required [liu2024adaptive]. Typical downstream tasks include sensitivity analysis [kihara2019estimating, ankenbrand2021sensitivity], gradient-based optimization [dogo2018comparative, daoud2023gradient], uncertainty quantification [abdar2021review, caldeira2020deeply, Zhong2023PIVAE], digital twins [chattopadhyay2024oceannet, kobayashi2024deep, kobayashi2024improved], and inverse or generative design [liu2025univariate, liu2025towards]. These surrogates do not replace a single high-fidelity simulation; rather, they amortize its upfront cost over large query workloads, enabling dramatically faster iterative engineering and design cycles.
@@
where ${G}_{\theta}$ is the neural operator with learnable parameter $\theta$, $\mathcal{F}$ is the space of input functions, and $\mathcal{G}$ is the space of output functions on query points x. Examples such as the Fourier Neural Operator (FNO) [li2020fourier] and the Deep Operator Network (DeepONet) [lu2021learning] have shown significant success in this area [kovachki2021universal, you2022learning, bonev2023spherical, cai2024towards, he2023novel, li2023phase, he2024sequential, gano, dcon, fc4no]. These models map inputs like initial and boundary conditions, material properties, and other functions to solution fields at query points. However, their applicability is often limited to simple geometries. For instance, FNO is tailored for fixed, regular meshes such as image grids or 3D voxels and requires a complete mesh graph as input due to its reliance on the Fast Fourier Transform kernel, making it unsuitable for arbitrary geometries and less scalable. While DeepONet can predict solution fields at arbitrary query points, its original formulation is restricted to single geometries with fixed query point locations and counts.
@@
However, real-world problems frequently involve complex and arbitrary computational geometries. Recent advances, including graph neural operators (GNO) [li2020neural], Geo-FNO [li2023fourier], GI-FNO [li2023GIFNO], and Geom-DeepONet [he2024geom], have addressed challenges in handling diverse geometries. GNO can process irregular grids using graph representations but is limited to local operations under constrained computational budgets, resulting in lower accuracy. Both Geo-FNO and GI-FNO rely on projecting irregular meshes onto fixed regular grids and mapping them back. However, Geo-FNO requires a fixed number of query points during training, while GI-FNO constructs new graphs on regular meshes using local kernel integration layers through GNO, which need longer inference time. Geom-DeepONet, while overcoming the limitation of single geometries of DeepONet by removing or repeating nodes to ensure fixed data size, is restricted to parametric geometries (e.g., length, thickness, radius) and does not generalize to fully arbitrary shapes. Additionally, GI-FNO requires a signed distance function (SDF) as input to represent geometry, and Geom-DeepONet depends on both geometry parameters and the SDF. However, obtaining the SDF is not always easily, particularly for complex 3D geometries represented by surface meshes, as it incurs a computational cost of $O(N_{g}\times N_{m})$ for distance calculation and $O(N_{g}\times N_{m})$ for the minimum value, where $N_{g}$ is the number of grid points in the domain and $N_{m}$ is the number of surface mesh—both of which are large for intricate geometries.
@@
Researchers have leveraged the attention mechanism of transformers [Ashish2017atten] in neural operators to efficiently integrate information from input functions and query points [cao2021choose, liu2022ht, li2022transformer, hao2023gnot]. Neural Operator Transformers (NOT) are particularly advantageous for irregular meshes due to their attention mechanism, enabling the prediction of solution fields at arbitrary query points. Unlike DeepONet, which employs a simple dot product to combine information from input functions and query points, the attention mechanism in NOT directs each query point to focus on the most relevant information from the input functions. This capability allows NOT to generalize effectively and capture more complex relationships between input functions and query points. Building on this, the authors developed an SDF-based NOT that accommodates arbitrary geometries [liu2025towards]. The SDF representation facilitates handling arbitrary geometries on fixed, regular grids, which are compatible with existing deep learning models. However, as noted earlier, obtaining the SDF is not always straightforward or computationally easily.
@@
In this work, we introduce the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries without requiring SDF as geometry features. GINOT consists of two primary components: the geometry encoder and the solution decoder. Drawing inspiration from Shap-E [jun2023shap] by OpenAI, we design the geometry encoder to process the boundary point cloud, representing the geometry, into the KEY and VALUE matrices for the attention mechanism. This choice facilitates the integration of information from query points in the solution decoder.
@@
The geometry encoder utilizes sampling and grouping layers [qi2017pointnet, qi2017pointnet++] to extract local geometric features from the boundary point cloud. These local features are fused with global geometry information through cross-attention layers, enabling effective processing of unordered and non-uniformly distributed point clouds—unlike convolutional neural networks (CNNs), which operate on regular grids with uniform density. Because the number of points in the point cloud can vary between geometries, padding is applied to standardize input sizes within a batch. To ensure that padding points do not influence the geometry encoder’s output, a masking mechanism is employed: it prevents padding points from being selected in the sampling and grouping layers and excludes them from contributing to attention scores in the cross-attention layer.
@@ 2 GINOT
GINOT integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries. It consists of two primary components: the geometry encoder and the solution decoder, as illustrated in Fig.˜1. Both components leverage the attention mechanism inherent in the transformer architecture [Ashish2017atten], which facilitates the model’s ability to focus on relevant input information. The attention mechanism is defined as
@@
where $Q\in\mathbb{R}^{N_{q}\times d_{e}}$ represents the QUERY matrix, and $K,V\in\mathbb{R}^{N_{k}\times d_{e}}$ are the KEY and VALUE matrices, respectively. Here, $d_{e}$ denotes the embedding dimension, $N_{q}$ is the length of the query sequence, and $N_{k}$ is the length of the KEY and VALUE sequences. To ensure stable training, residual connections and layer normalization are applied within each attention block, as depicted in the bottom right of Fig.˜1.
@@ 2.1 Geometry encoder
To address these challenges, we developed a geometry encoder that processes the boundary point cloud of a geometry into the KEY and VALUE matrices for the attention mechanism. This enables seamless integration of geometry information with query points in the solution decoder. In contrast to CNN inputs, boundary point cloud used to represent geometries have three key characteristics: (1) invariance to the order of points, (2) non-uniform point density across locations, and (3) a variable number of points across geometries. Drawing inspiration from Shap-E [jun2023shap] by OpenAI, the geometry encoder, as shown in Fig.˜1(a), is designed to handle these characteristics. It ensures invariance to point order and padding while maintaining robustness to variations in point density.
@@
The point cloud is first processed through a sampling layer and a grouping layer for local geometry features, originally introduced in PointNet [qi2017pointnet] and PointNet++ [qi2017pointnet++], which are well-suited to handle the aforementioned properties. The sampling layer employs the iterative farthest point sampling (FPS) method [qi2017pointnet++] to select $N_{s}$ points from the point cloud. This method begins by randomly selecting a point, then iteratively selects the farthest point from the already selected points until $N_{s}$ points are chosen. Compared to random sampling, the FPS method provides better coverage of the entire point set for a given number of centroids. The selected points are then passed to the grouping layer, which forms $N_{s}$ groups, each containing $N_{p}$ points.
@@
For each group, the $N_{p}$ points are selected as those within a ball of radius $r$ (a hyperparameter) centered at the corresponding sampled point. If fewer than $N_{p}$ points exist within the ball, the group is padded with the nearest points to the center. Conversely, if more than $N_{p}$ points are present, the nearest $N_{p}$ points are chosen. Consequently, the output of the grouping layer has dimensions $N_{s}\times N_{p}\times d$, where $d$ represents the dimensionality of the point cloud. The sampling and grouping layers allow to extract local geometry features from nonuniform density point cloud [qi2017pointnet++]. These layers output the selected points along with their indices in the original point cloud, which are subsequently used in downstream layers.
@@
The point cloud is also processed through a positional encoder layer utilizing the Nerf encoding method [mildenhall2021nerf], followed by a linear layer. Using the indices from the sampling and grouping layers, the output is indexed to extract local geometry information, resulting in a tensor of shape $N_{s}\times N_{p}\times C$, where $C$ represents the output channel dimension of the positional encoder. This tensor is concatenated with the output of the grouping layer. The combined tensor is then passed through a series multi-layer perceptron (MLP) layers, which aggregates the information within each group, reducing the dimension of $N_{p}$ to 1. As a result, the final output has a shape of $N_{s}\times C$.
@@
where the output is unaffected by permutations of the index $m$, which is a dummy index under Einstein summation notation. To ensure strict invariance, the sampling and grouping layers must also be invariant to the permutation of the point cloud. This requires consistently selecting the same initial point in the iterative FPS sampling layer and grouping points based on their distance to the centroids chosen by the sampling layer. However, as the sampling and grouping layers can sufficiently extract the geometry features regardless of the initial point selection, we use random initialization of the first point during FPS iteration in training. During inference, the first point is fixed as the first non-padding point of the point cloud to ensure consistent results. Our experiments demonstrate that this random initialization has a negligible impact on GINOT’s output (as shown in Section˜3.5).
The computational cost of the sampling and grouping layers is mainly attributed to the FPS and ball query operations. The FPS step, which iteratively selects the farthest point by computing distances between all pairs of points in the point cloud, has a complexity of $O(N^{2})$. The ball query step, which identifies points within a specified radius around each sampled point, involves a complexity of $O(N\times N_{s})$, where $N$ is the total number of points in the point cloud and $N_{s}$ is the number of sampled points. Importantly, both $N$ and $N_{s}$ are typically much smaller than the number of grid points and surface mesh points required in SDF-based methods, making the sampling and grouping layers in GINOT computationally more efficient, as further elaborated in Section˜3.6.
The sampling and grouping layers, combined with the attention mechanism, allow that the geometry encoder handle the point cloud with varying points density, while remaining robust to the permutation of points order. Since the number of points in the point cloud varies across different geometries, padding is typically required to standardize input sizes within a batch for deep learning models. To prevent padding points from influencing the geometry encoder’s output, a masking mechanism is employed. This mechanism excludes padding points from being selected in the sampling and grouping layers. Additionally, padding points can interfere with attention scores in the cross-attention layer of the geometry encoder. To mitigate this, the corresponding entries in the attention score matrix $(Q\cdot K^{T})/\sqrt{d_{e}}$ are set to $\infty"$, ensuring that the attention layer’s output is unaffected by padding points. The flexibility and robustness of the proposed geometry encoder are illustrated in experiments Section˜3.
@@ 2.2 Solution decoder
The solution decoder predicts the solution field at the query points using the geometry information from the geometry encoder. First, the query points are processed through a Nerf positional encoding, followed by an MLP, to generate the QUERY matrix with dimensions $N_{q}\times d_{e}$, where $N_{q}$ is the number of query points and $d_{e}$ is the embedding dimension. This QUERY matrix is then passed into cross-attention layers, which integrate the geometry information from the geometry encoder. The attention mechanism ensures that the query points focus on the most relevant geometry features. Finally, the output of the cross-attention layers is decoded into the solution field at the query points using another MLP. The architecture of the solution decoder is illustrated in Fig.˜1(b).
@@ 2.3 Extending GINOT
Many problems involve not only varying geometries but also additional input parameters, such as loading conditions, material properties, or boundary conditions. To address such cases, GINOT can be extended by incorporating additional encoders to process these input functions. The outputs of these additional encoders are concatenated with the output of the geometry encoder and passed through an MLP for aggregation of all information from the geometries and the additional inputs. The aggregated output is then used as the KEY and VALUE in the cross-attention blocks of the solution decoder. The architecture of the extended GINOT is illustrated in Fig.˜2. This extension is demonstrated in the bracket lug example in Section˜3.4, where varying loading conditions are included as additional inputs.
@@ 3 Experiments
where $y^{\text{true}}$ and $y^{\text{pred}}$ represent the ground truth and predicted solutions, respectively.
@@ 3.1 Overview
To evaluate the performance and accuracy of GINOT, six datasets are utilized: elasticity [li2023fourier], Poisson equation with both structured and unstructured meshes, bracket lugs [he2024geom], micro-periodic unit cell (PUC) [liu2025dataset] in [liu2025towards], and Jet Engine Bracket (JEB) [hong2024deepjeb]. Table˜1 provides a detailed summary of these datasets, including their size, the number of query points, the number of points in the boundary point cloud, whether the dataset is parametric or fully arbitrary, and the type of solution considered. The first three datasets have a constant number of points in the boundary point cloud, while the last three datasets feature a varying number of points in the boundary point cloud. All datasets feature non-uniform point densities, as illustrated in the visualization results in the following subsections. While the datasets considered here focus on representative linear and nonlinear single-physics cases, the framework is not confined to single-physics or weakly coupled settings. With suitably generated training data, it works for complex, strongly coupled multi-physics problems.
@@
DataSet Elasticity[li2023fourier]
Poisson
(structure mesh)
Poisson
(unstructure mesh)
Bracket lug [he2024geom]
Micro-
PUC [liu2025dataset]
JEB [hong2024deepjeb]
@@
Number of
query point
972 2737 [1566,2838] [5054,15210] [1136,$63$] [18085,55009]
Number points
in boundary PC
105 144 144 [2184,4284] [180,484] [2253,10388]
@@
Solution Stress Mises stress
Mises stress
displacements
Mises stress
Table˜2 summarizes the training hyperparameters of the GINOT models. For the micro PUC dataset, two models are trained: one for predicting the solution at the final strain step and another for predicting the solution over historical strain steps.
@@
Elasticity
Poisson
(structure
mesh)
Poisson
(unstructure
mesh)
Bracket
lug
Micro PUC
(Final
strain step)
Micro PUC
(historical
strain steps)
JEB
Batch size 20 32 32 32 64 64 16
Optimizer Adam Adam Adam Adam Adam Adam AdamW
Initial learning rate 0.001 0.001 0.001 0.001 0.001 0.001 0.0005
@@
Grouping $r$ 0.2 0.2 0.2 0.5 0.1 0.2 0.1
Att. heads (Decoder) 4 8 8 8 4 4 4
@@
Table˜3 summarizes the computational efficiency of GINOT. Training tasks were conducted on single NVIDIA A100 GPU on the DELTA machine or single NVIDIA H100 GPU on the DeltaAI machine, both hosted at the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign (UIUC). Inference tasks were executed on single NVIDIA A100 GPU on the DELTA machine.
@@
Elasticity
Poisson
(structure
mesh)
Poisson
(unstructure
mesh)
Bracket
lug
Micro PUC
(Final
strain step)
Micro PUC
(historical
strain steps)
JEB
@@
Table˜4 summarizes GINOT’s performance across the datasets, presenting the $L_{2}$ relative error for both training and testing samples. These results are compared against baseline methods, with the best performance highlighted in bold. The findings demonstrate GINOT’s ability to tackle more complex problems while outperforming the baseline methods. Detailed experimental results are provided in the subsequent subsections.
@@
Dataset, Metric Data Geo-FNO[li2023fourier]
Geom-
DeepONet[he2024geom]
GNO[li2020neural] GI-FNO[li2023GIFNO] GINOT
Elasticity, $L_{2}$ Training 1.25% NA 11.3% 1.32% 0.806%
Testing 2.29% NA 12.3% 1.66% 1.03%
Poisson, $L_{2}$
(structure mesh)
Training 0.43% NA 7.43% 0.42% 0.39%
@@
Poisson, $L_{2}$
(unstructure mesh)
Training NA 7.98% 0.49% 0.41%
@@
[…]
Testing 1.824% 11.3% 2.81% 0.40%
Micro PUC, $L_{2}$
(Final strain step)
Training 16.5% 9.24% 6.99%
Testing 18.2% 9.95% 7.61%
Micro PUC, $L_{2}$
(historical strain steps)
Training 18.9% 9.56% 8.42%